Estimate the free energy change (ΔGrxn) for the following reaction at 298 K, if each gaseous substance is present at p = 0.70 atm. C2H5OH(l) + 3O2(g) ⇌ 2CO2(g) + 3H2O(l)
Answer #1

Cq H₃ 04 (d) + 30g (9) 210g (g) + 342010) k = [cog] - (0.70² - 1 = 1.43 (0.703 0.70 Togli? Ah= - Rilak E-8.314x298 x In (143)